OpenStudy (anonymous):

One more integrate y^2 sqrt(y)dy

OpenStudy (zarkon):

\[y^2\sqrt{y}=y^2y^{1/2}=y^{2+(1/2)}=y^{5/2}\] now integrate

OpenStudy (anonymous):

oh okay awesome thank you